OpenStudy (anonymous):

Estimate the area of a circle with a radius of 16 meters. Use 22/7 for π.

OpenStudy (anonymous):

62 square meters 531 square meters 805 square meters 1,608 square meters

OpenStudy (anonymous):

@rebeccaxhawaii @KendrickLamar2014 @welshfella

OpenStudy (anonymous):

do you know how to find the area of a circle?

OpenStudy (anonymous):

do you know how to find the area of a circle?

OpenStudy (anonymous):

I cant remember @TwentyOnePilots_

OpenStudy (palwandew):



OpenStudy (anonymous):

@haleyelizabeth2017 please help me?

OpenStudy (anonymous):

so you set up your equation we konwthat the radius is 16 so we just plug it in A= 3.14 * 16 *2

OpenStudy (anonymous):

100.48 @TwentyOnePilots_

OpenStudy (welshfella):

(22/7) * 16^2

OpenStudy (anonymous):

actually i think your supposed to times the radius by its self so it would be 3.14*16*16 which then equals 804.25

OpenStudy (anonymous):

Thats not an answer choice though @TwentyOnePilots_

OpenStudy (anonymous):

sorry my calculations were wrong 3.14 (exactly) * 16*16 =803.84 which when rounded is 804 square meters

OpenStudy (anonymous):

which still isnta answer :( sorry i tried my best let me call some one @imqwerty

imqwerty (imqwerty):

well they asked you to use \(\large \frac{22}{7}\) as \(\pi\) @TwentyOnePilots_ used \(\pi=3.14\) which is alright but we must do as the question says so we will take \(\large \pi=\frac{22}{7}\) :) now we use this equation to find the area-> \(\large Area= \pi r^2\) \(\large Area= \frac{22}{7} \times 16^2\) well the Area that you will be getting from this equation will be in decimal form and you need to round it to the nearest integer to get your solution (:

OpenStudy (anonymous):

509.14 is what I got @imqwerty

imqwerty (imqwerty):

um check your calculations again i'd simplify the thing a bit- \(\large Area= \frac{22}{7} \times 16^2\) \(\large Area= \frac{22}{7} \times (16 \times 16)\) \(\large Area= \frac{22}{7} \times 256\) \(\large Area= \frac{5632}{7} \)

OpenStudy (anonymous):

I cant figure it out @imqwerty

imqwerty (imqwerty):

are those steps clear ? is the problem coming in the calculation part?

OpenStudy (anonymous):

I thinks it 531 @imqwerty

imqwerty (imqwerty):

\(\large Area= \frac{5632}{7}\)